# Accurate and Efficient Low-Rank Model Merging in Core Space

### v1

@@ 5 Experimental Results
Baseline merging methods. We evaluate each merging space using the following merging methods. Task Arithmetic (TA) (ilharco2023task, ) performs a scaled summation of each task matrix $W_{\text{merged}}=W_{0}+\alpha\sum_{i=1}^{T}\Delta W_{i}$. As this is a linear operation, the results of merging in each space are the same (see Eq.˜13 for Core and stoica2024knots for KnOTS). TIES (yadav2023tiesmerging, ) trims low-magnitude parameters and averages parameters with dominating sign, while DARE (yu2024language, ) preprocesses task vectors by randomly dropping a fraction of parameters and rescaling the remaining ones. TSV (tsv, ) concatenates low-rank approximations of task matrices and orthogonalizes them across tasks. CART choi2024revisiting calculates centered task vectors as a difference of fine-tuned weights from the average of all fine-tuned weights and performs task arithmetic on the low-rank approximation of these centered task vectors. Iso-C (marczak2025task, ) flattens the spectrum of singular values for a model merged with task arithmetic. As the spectrum flattening can be performed on weights merged with any merging technique, we combine Iso with other merging techniques, denoting it with +Iso-C.
@@ 5.1 Results
Joint-task evaluation in vision setting. We also evaluate vision models in the challenging joint-task setting introduced in stoica2024knots , in which the task ID is unknown during inference. Instead of performing a multi-task evaluation, it evaluates the merged model based on the union of all classes, requiring the model to distinguish between classes from all tasks. We present the results in Tab.˜4. Core Space facilitates merging with almost all methods, achieving state-of-the-art results when combined with Iso-C.
@@ 5.2 Analysis
Truncation. In this Section, we compare the utilization of full subspace and core space for models merged with TA. Firstly, we calculate the SVD of the merged matrices: $\Delta W_{\text{merged}}$ for full space and $M_{\text{merged}}$ for core space. Then, we truncate a fraction of $p$ least significant values, i.e., $\sigma_{i}=0$ for $i>(1-p)*\text{dim}(\Sigma)$, and observe the drop in the performance of the merged model after truncation. We present the results in Fig.˜5. In full space, we can truncate a fraction up to $p=0.8$ values without performance loss, while in core space, truncation of any component results in a performance drop. It shows that the core space is dense while the full space contains many unused or redundant components. We hypothesize that the compactness of core space facilitates model merging as it extracts only the relevant components.
@@
Table 5: Ablations on the choice of reference basis. Our selected basis (3) achieves significantly higher results than the single-task basis (1) and the random orthonormal basis of the same dimensionality (2). We proceed with $V_{A}^{\mathrm{ref}}$ analogously to $U_{B}^{\mathrm{ref}}$. We report the TIES-Core results on vision tasks with ViT-B/32.
@@
[](3) Concatenation (Eq.˜4) $m\times Tr$ 68.6 0.0
To evaluate the reference bases choice, we compare the performance of merging in core space against different reference bases, and compute the alignment error $\varepsilon_{U}$ defined in Eq.˜10 (averaged over all layers and tasks). We present the results in Tab.˜5. In row (1), we evaluate using the basis of the first task as a reference basis. In row (2), we set the reference basis to a random orthonormal basis of the same dimensionality. These two bases perform much less than our reference basis in row (3). Moreover, we confirm that the optimal reference basis from row (3) achieves zero alignment error. Additionally, we verified experimentally that even in the extreme case where $T\cdot r>\min(m,n)$ (e.g., $Tr=2048>768$ for merging 8 ViT-B/32 LoRA models), the reconstruction error defined in Eq.˜10 remains exactly zero, consistent with the generalized theoretical result in Sec.˜A.4.
@@ 6 Conclusion
We propose Core Space, an efficient and effective method for merging task-specific LoRA modules. By projecting task-specific LoRA updates into a common subspace, Core Space reduces alignment error, leading to consistent accuracy improvements and SOTA results in both vision and language settings, while remaining computationally efficient. Our evaluations across vision and language domains confirm its scalability and strong performance in practical settings. We believe that Core Space can contribute to more efficient and accessible model adaptation in multi-task settings, particularly for large models.
@@ Acknowledgments
We acknowledge the Spanish project PID2022-143257NB-I00, financed by MCIN/AEI/10.13039/501100011033 and FEDER, and Funded by the European Union ELLIOT project. This work was supported by the Fortissimo Plus (FFplus) project Grant Agreement No. 101163317), under the European High‑Performance Computing Joint Undertaking (EuroHPC JU) and the Digital Europe Programme. The authors gratefully acknowledge access to compute resources enabled by FFplus. Funded by the European Union. Aniello Panariello acknowledges travel support from ELIAS (GA no 101120237). Daniel Marczak is supported by National Centre of Science (NCN, Poland) Grant No. 2021/43/O/ST6/02482. Bartłomiej Twardowski acknowledges the grant RYC2021-032765-I and National Centre of Science (NCN, Poland) Grant No. 2023/51/D/ST6/02846.
@@ A.1 Least Squares Solutions
Consider the problems in Eq.˜14. Each objective is convex and admits a unique global minimizer, as $U_{B}^{\mathrm{ref}}$ and $V_{A}^{\mathrm{ref}}$ have full column rank due to their orthonormality. To solve the first problem, we compute the gradient of the objective function with respect to $R$:
@@
Similarly, for the second problem in Eq.˜14:
@@ A.2 Alignment Error Quantification
To derive the alignment error, we simplify the notation by temporarily omitting the dependency on the model index $t$. Substituting the definition $R_{B}={U_{B}^{\mathrm{ref}}}^{\top}U_{B}$ from Eq.˜15, we can write:
@@ A.3 Optimal Reference Bases
is given by any orthonormal basis whose columns include the top $r$ eigenvectors corresponding to nonzero eigenvalues of $B^{(t)}{B^{(t)}}^{\top}\in\mathbb{R}^{m\times m}$ or, equivalently, by the top $r$ left singular vectors of $B^{(t)}$. At the optimum, the objective achieves the maximum value of $r$, yielding zero alignment error in $\varepsilon_{U}$ as defined in Eq.˜16.
@@
The second condition, Eq.˜21, holds by construction: we explicitly choose $U_{*}\in\mathbb{R}^{m\times T\cdot r}$ to be an orthonormal matrix. Specifically, we define $U_{*}=[v_{1},\ldots,v_{r},p_{1},\ldots,p_{(T-1)\cdot r}]$, where $\{v_{i}\}_{i=1}^{r}$ are orthonormal eigenvectors of $U_{B}^{(t)}U_{B}^{(t)\top}$ associated with its nonzero eigenvalues $\lambda_{1},\ldots,\lambda_{r}$, and $\{p_{j}\}_{j=1}^{(T-1)\cdot r}$ are additional orthonormal vectors chosen to complete the basis.
Next, to verify the first condition reported in Eq.˜20, we define the diagonal matrix $\Lambda_{*}=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{r},0,\ldots,0)\in\mathbb{R}^{T\cdot r\times T\cdot r}$, where the trailing zeros correspond to the eigenvalues associated with the orthogonal complement. Since each $v_{i}$ is an eigenvector of $U_{B}^{(t)}U_{B}^{(t)\top}$ with eigenvalue $\lambda_{i}$, and each $p_{j}$ lies in the nullspace of that matrix, we have:
@@
which confirms that the first-order condition in Eq.˜20 is satisfied.
@@ A.4 Generalization Beyond the $T\cdot r\leq m,n$ Assumption
Least-Squares Solution. Rewriting the least-squares problem (Eq.˜14) in terms of the truncated $U_{B}^{\text{ref}}$ gives:
@@ Appendix D Additional Experiment Details
D.1 Optimized Pipeline
Figure 7: The proposed Core Space Merging with the optimized algorithm. First, we perform SVD on a concatenation of low-dimensional $A^{(t)}$ and $B^{(t)}$ matrices to obtain reference bases. Afterwards, we directly compute the aligned core matrices. Finally, we perform merging in the Core Space and reconstruct to obtain the final $\Delta W$.
We report in Fig.˜7 the illustration of the Core Space approach with the optimized pipeline presented in Sec.˜4.2.
D.2 Experimental Environment
@@ D.3 Hyperparameter Search
D.4 Code Availability and Pseudocode
@@
Fig.˜8 gives PyTorch-style pseudocode illustrating how to apply any merging strategy within the Core Space framework.
@@
# Theory Version
# U_A, S_A, Vh_A = svd(A, full_matrices=False) # (r, r), (r,), (r, n)
# U_B, S_B, Vh_B = svd(B, full_matrices=False) # (m, r), (r,), (r, r)
[…]
# Q_A = Vh_A @ Vh_A_ref.T
# R_B = U_B_ref.T @ U_B
[…]
# M_aligned = R_B @ M @ Q_A # (T*r, T*r)
[…]
M_aligned = U_B_ref.T @ B @ A @ Vh_A_ref.T

### v3

@@ 5 Experimental Results
Baseline merging methods. We evaluate each merging space using the following merging methods. Task Arithmetic (TA) (ilharco2023task, ) performs a scaled summation of each task matrix $W_{\text{merged}}=W_{0}+\alpha\sum_{i=1}^{T}\Delta W_{i}$. As this is a linear operation, the results of merging in each space are the same (see Eq.˜11 for Core and stoica2024knots for KnOTS). TIES (yadav2023tiesmerging, ) trims low-magnitude parameters and averages parameters with dominating sign, while DARE (yu2024language, ) preprocesses task vectors by randomly dropping a fraction of parameters and rescaling the remaining ones. TSV (tsv, ) concatenates low-rank approximations of task matrices and orthogonalizes them across tasks. CART choi2024revisiting calculates centered task vectors as a difference of fine-tuned weights from the average of all fine-tuned weights and performs task arithmetic on the low-rank approximation of these centered task vectors. Iso-C (marczak2025task, ) flattens the spectrum of singular values for a model merged with task arithmetic. As the spectrum flattening can be performed on weights merged with any merging technique, we combine Iso with other merging techniques, denoting it with +Iso-C.
@@ 5.1 Results
Joint-task evaluation in vision setting. We also evaluate vision models in the challenging joint-task setting introduced in stoica2024knots , in which the task ID is unknown during inference. Instead of performing a multi-task evaluation, it evaluates the merged model on the union of all classes, requiring the model to distinguish between classes from all tasks. We present the results in Tab.˜4. Core Space facilitates merging with almost all methods, achieving state-of-the-art results when combined with Iso-C.
@@ 5.2 Analysis
Truncation. We herein compare the utilization of full space and Core Space for models merged with TA. Firstly, we calculate the SVD of the merged matrices: $\Delta W_{\text{merged}}$ for full space and $M_{\text{merged}}$ for Core Space. Then, we truncate a fraction of $p$ least significant values, i.e., $\sigma_{i}=0$ for $i>(1-p)*\text{dim}(\Sigma)$, and observe a drop in the accuracy of the merged model after truncation. As shown in Fig.˜5, in full space we can truncate a fraction up to $p=0.8$ values without performance loss, while in core space, truncation of any component results in a performance drop. It shows that the core space is dense while the full space contains many unused or redundant components. We hypothesize that the compactness of core space facilitates model merging as it extracts only the relevant components.
@@
Table 5: Ablations on the choice of reference basis. Our basis (3) achieves higher results than the single-task basis (1) and the random orthonormal basis of the same dimensionality (2). We proceed with $V_{A}^{\mathrm{ref}}$ analogously to $U_{B}^{\mathrm{ref}}$. We report the TIES-Core results on vision tasks with ViT-B/32.
@@
[](3) Concatenation (Eq.˜3) $m\times Tr$ 68.6 0.0
To evaluate the reference bases choice, we assess the performance when adopting different reference bases, and compute the alignment error $\varepsilon_{U}$ defined in Eq.˜13 (averaged over all layers and tasks). We present the results in Tab.˜5. In row (1), we evaluate using the basis of the first task as a reference basis. In row (2), we set the reference basis to a random orthonormal basis of the same dimensionality. These two bases perform much less than our reference basis in row (3). Moreover, we confirm that the optimal reference basis from row (3) achieves zero alignment error. Additionally, we verified experimentally that even in the extreme case where $T\cdot r>\min(m,n)$ (e.g., $Tr=2048>768$ for merging 8 ViT-B/32 LoRA models), the reconstruction error defined in Eq.˜13 remains exactly zero, consistent with the generalized theoretical result in Sec.˜A.4.
@@ 6 Conclusion
We propose Core Space, an efficient and effective method for merging LoRA modules. By projecting task-specific LoRA updates into a common subspace, Core Space reduces alignment error, leading to consistent accuracy improvements and SOTA results in both vision and language settings, while remaining computationally efficient. Our evaluations across vision and language domains confirm its scalability and strong performance in practical settings. We believe that Core Space can contribute to more efficient and accessible model adaptation in multi-task settings, particularly for large models.
@@ Acknowledgments
We acknowledge the Spanish project PID2022-143257NB-I00, financed by MCIN/AEI/10.13039/501100011033 and ERDF/EU, and Funded by the European Union ELLIOT project. This work was supported by the Fortissimo Plus (FFplus) project Grant Agreement No. 101163317), under the European High‑Performance Computing Joint Undertaking (EuroHPC JU) and the Digital Europe Programme. The authors gratefully acknowledge access to compute resources enabled by FFplus. Funded by the European Union. Aniello Panariello acknowledges travel support from ELIAS (GA no 101120237). Daniel Marczak is supported by National Centre of Science (NCN, Poland) Grant No. 2021/43/O/ST6/02482. Bartłomiej Twardowski acknowledges the grant RYC2021-032765-I and National Centre of Science (NCN, Poland) Grant No. 2023/51/D/ST6/02846.
@@ A.1 Least Squares Solutions
Consider the problems in Eq.˜16. Each objective is convex and admits a unique global minimizer, as $U_{B}^{\mathrm{ref}}$ and $V_{A}^{\mathrm{ref}}$ have full column rank due to their orthonormality. To solve the first problem, we compute the gradient of the objective function with respect to $R$:
@@
Similarly, for the second problem in Eq.˜16:
@@ A.2 Alignment Error Quantification
To derive the alignment error, we simplify the notation by temporarily omitting the dependency on the model index $t$. Substituting the definition $R_{B}={U_{B}^{\mathrm{ref}}}^{\top}U_{B}$ from Eq.˜17, we can write:
@@ A.3 Optimal Reference Bases
is given by any orthonormal basis whose columns include the top $r$ eigenvectors corresponding to nonzero eigenvalues of $B^{(t)}{B^{(t)}}^{\top}\in\mathbb{R}^{m\times m}$ or, equivalently, by the top $r$ left singular vectors of $B^{(t)}$. At the optimum, the objective achieves the maximum value of $r$, yielding zero alignment error in $\varepsilon_{U}$ as defined in Eq.˜18.
@@
The second condition, Eq.˜23, holds by construction: we explicitly choose $U_{*}\in\mathbb{R}^{m\times T\cdot r}$ to be an orthonormal matrix. Specifically, we define $U_{*}=[v_{1},\ldots,v_{r},p_{1},\ldots,p_{(T-1)\cdot r}]$, where $\{v_{i}\}_{i=1}^{r}$ are orthonormal eigenvectors of $U_{B}^{(t)}U_{B}^{(t)\top}$ associated with its nonzero eigenvalues $\lambda_{1},\ldots,\lambda_{r}$, and $\{p_{j}\}_{j=1}^{(T-1)\cdot r}$ are additional orthonormal vectors chosen to complete the basis.
Next, to verify the first condition reported in Eq.˜22, we define the diagonal matrix $\Lambda_{*}=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{r},0,\ldots,0)\in\mathbb{R}^{T\cdot r\times T\cdot r}$, where the trailing zeros correspond to the eigenvalues associated with the orthogonal complement. Since each $v_{i}$ is an eigenvector of $U_{B}^{(t)}U_{B}^{(t)\top}$ with eigenvalue $\lambda_{i}$, and each $p_{j}$ lies in the nullspace of that matrix, we have:
@@
which confirms that the first-order condition in Eq.˜22 is satisfied.
@@ A.4 Generalization Beyond the $T\cdot r\leq m,n$ Assumption
Least-Squares Solution. Rewriting the least-squares problem (Eq.˜16) in terms of the truncated $U_{B}^{\text{ref}}$ gives:
@@ Appendix D Additional Experiment Details
D.1 Experimental Environment
@@ D.2 Hyperparameter Search
D.3 Code Availability and Pseudocode
@@
Fig.˜7 gives PyTorch-style pseudocode illustrating how to apply any merging strategy within the Core Space framework.
@@
M_aligned = (U_B_ref.T @ B) @ (A @ Vh_A_ref.T)